OpenStudy (anonymous):

a gardener is planning to make a rectangular garden with an area of 80ft^2.she has 12 yards of fencing to put around the perimeter of the garden.what should the dimension of the garden be ? please answer :))

OpenStudy (anonymous):

please answer !! i need your help

OpenStudy (anonymous):

you put yards into feets so units are the same and you can solve the problem by this equation system: 2a+2b=12 (put this number in feets) a*b=80

OpenStudy (anonymous):

huh ? Ivcvs , explain it further .. please

OpenStudy (anonymous):

you solve the equation system for a and b, and you have the dimensions of the garden

OpenStudy (anonymous):

a=length, b=width

OpenStudy (anonymous):

so what is a and what is b ?

OpenStudy (anonymous):

length yung a yung b width

OpenStudy (anonymous):

pano mo nga makkha un ?

OpenStudy (anonymous):

ung a at b ?

OpenStudy (anonymous):

w8

OpenStudy (anonymous):

?? xd

OpenStudy (anonymous):

how to solve the a and the b

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

by the way ty lvcvs

OpenStudy (anonymous):

a=80/b; so we subsitute in first equation and we have; 2*(80/b)+2b=12; 160/b+2b=12; 160+2b^2/b=12; 160+2b^2=12b; 2b^2-12b+160=0; here's a quadratic equation, do you know how to solve it?

OpenStudy (anonymous):

hahahaha its difficult

OpenStudy (anonymous):

we havent discussed it

OpenStudy (anonymous):

can you please dicuss it to us ?? please ??

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

excited ka pando

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

why is there 2 b ?

OpenStudy (anonymous):

let the to sides be x nd y, so arey=x*y=80 and x+y=12/2=6 yards, now convert it into feets= 18 feets so x=18-y puttting in 1st equation, y(18-y)=80 clearly the asnwer is 8 nd 10